# Constructing Concise Characteristic Samples for Acceptors of Omega Regular Languages

### v3

@@ 1. Introduction
One of the main obstacles in obtaining a polynomial learning algorithm for regular $\omega$-languages is that they do not in general have a Myhill-Nerode characterization; that is, there is no theorem correlating the states of a minimal automaton of some of the common automata types (Büchi, parity, Muller, etc.) to the equivalence classes of the right congruence of the language. The right congruence relation for an $\omega$-language $L$ relates two finite words $x$ and $y$ iff there is no infinite suffix $z$ differentiating them, that is $x\sim_{L}y$ (for $x,y\in\Sigma^{*}$) iff $\forall z\in\Sigma^{\omega}.\ xz\in L\iff yz\in L$. The quest for finding a polynomial query learning algorithm for a subclass of the regular $\omega$-languages, led to studying subclasses of languages for which such a relation holds. These languages are termed fully informative [AF18]. We use $\mathbb{IBA},\mathbb{ICA},\mathbb{IPA},\mathbb{IRA},\mathbb{ISA},\mathbb{IMA}$ to denote the classes of languages that are fully informative of type Büchi, coBüchi, parity, Rabin, Streett, and Muller, respectively. A language $L$ is said to be fully informative of type $\mathbb{X}$ for $\mathbb{X}\in\{\mathbb{B},\mathbb{C},\mathbb{P},\mathbb{R},\mathbb{S},\mathbb{M}\}$ if there exists a deterministic automaton of type $\mathbb{X}$ that recognizes $L$ and is isomorphic to the automaton derived from $\sim_{L}$. While many properties of these classes are now known, in particular that they span the entire hierarchy of regular $\omega$-languages [Wag75], a polynomial learning algorithm for them is not known.
We show (in Sections 5-9) that the classes $\mathbb{IBA},\mathbb{ICA},\mathbb{IPA},\mathbb{IRA},\mathbb{ISA},\mathbb{IMA}$ can be identified in the limit using polynomial time and data. We further show (in Section 10) that there is a polynomial time algorithm to compute a characteristic sample given an acceptor $\mathcal{C}\in\mathbb{IXA}$. To show that these classes are also efficiently teachable we need polynomial time algorithms for inclusion and equivalence of automata of these types, that also return shortlex counterexamples in case of inequivalence. Such an algorithm is known to exist for the classes $\mathbb{NBA}$, $\mathbb{NCA}$, $\mathbb{NPA}$, since these classes have inclusion algorithms in NL [Sch10]. For the other classes a polynomial-time algorithm can be obtained following a reduction to model checking a certain fragment of CTL formulas [CDK93]. However this reduction does not yield shortlex counterexamples. [^3] The formal definition of shortlex is deferred to Section 2. We provide such algorithms in Sections 11-14.
@@ 2. Preliminaries
Let ${{\mathcal{M}}}_{1}$ and ${{\mathcal{M}}}_{2}$ be two deterministic complete automata with the same alphabet $\Sigma$, where for $i=1,2$, ${{\mathcal{M}}}_{i}=\langle\Sigma,Q_{i},(q_{\iota})_{i},\delta_{i}\rangle$. Their product automaton, denoted ${{\mathcal{M}}}_{1}\times{{\mathcal{M}}}_{2}$, is the deterministic complete automaton ${\mathcal{M}}=\langle\Sigma,Q,q_{\iota},\delta\rangle$ such that $Q=Q_{1}\times Q_{2}$ is the set of ordered pairs of states of ${{\mathcal{M}}}_{1}$ and ${{\mathcal{M}}}_{2}$; the initial state $q_{\iota}=((q_{\iota})_{1},(q_{\iota})_{2})$ is the pair of initial states of the two automata; and for all $(q_{1},q_{2})\in Q$ and $\sigma\in\Sigma$, $\delta((q_{1},q_{2}),\sigma)=(\delta_{1}(q_{1},\sigma),\delta_{2}(q_{2},\sigma))$. For $i=1,2$, let $\pi_{i}$ be projection onto the $i$-th coordinate, so that for a subset $S$ of $Q$, $\pi_{1}(S)=\{q_{1}\in Q_{1}\mid\exists q_{2}\in Q_{2}(q_{1},q_{2})\in S\}$, and analogously for $\pi_{2}$.
@@
For finite words the acceptance condition is a set $F\subseteq Q$ and a run on a word $v$ is accepting if it ends in an accepting state, i.e., if $\delta(v)$ contains an element of $F$. For infinite words, there are various acceptance conditions in the literature, and we consider six of them: Büchi, coBüchi, parity, Rabin, Streett, and Muller, all based on the set of states visited infinitely often in a given run. For each model we define the related quantity of the size of the acceptor, taking into account the acceptance condition.
@@
We use three-letter acronyms for automata and classes concerning the common $\omega$-automata discussed above. The first letter is in $\{N,D,I\}$ and stands for non-deterministic, deterministic and isomorphic (or fully informative) which will be explained in the sequel. The second letter describes the acceptance condition, and the third letter $A$ stands for acceptor. That is, we use NBA, NCA, NPA, NRA, NSA, NMA (resp., DBA, DCA, DPA, DRA, DSA, DMA) for non-deterministic (resp., deterministic) Büchi, coBüchi, parity, Rabin, Street, and Muller acceptors. We use blackboard font for the respective classes of representations. That is, we use $\mathbb{NBA}$, $\mathbb{NCA}$, $\mathbb{NPA}$, $\mathbb{NMA}$, $\mathbb{NRA}$, and $\mathbb{NSA}$ (resp., $\mathbb{DBA}$, $\mathbb{DCA}$, $\mathbb{DPA}$, $\mathbb{DRA}$, $\mathbb{DSA}$, and $\mathbb{DMA}$) for the corresponding class of representations. It is known that $\mathbb{NCA}$ and $\mathbb{DCA}$ recognize the same languages and that the classes $\mathbb{DCA}$ and $\mathbb{DBA}$ are distinct proper subclasses of the regular $\omega$-languages. The other classes are the full class of regular $\omega$-languages.
@@
    Let $\mathcal{B}=\langle\Sigma,Q,q_{\iota},\delta,F\rangle$, where $\mathcal{B}$ is an NBA. Define the NRA $\mathcal{R}=\langle\Sigma,Q,q_{\iota},\delta,\{(F,\emptyset)\}\rangle$. The $\mathcal{B}$ and $\mathcal{R}$ are equivalent and have the same size. $\mathcal{R}$ is deterministic if $\mathcal{B}$ is.
@@
    Let $\mathcal{C}=\langle\Sigma,Q,q_{\iota},\delta,F\rangle$, where $\mathcal{C}$ is a NCA. Define the NSA $\mathcal{S}=\langle\Sigma,Q,q_{\iota},\delta,\{(F,\emptyset)\}\rangle$. Then $\mathcal{S}$ and $\mathcal{C}$ are equivalent and have the same size. $\mathcal{S}$ is deterministic if $\mathcal{C}$ is.
@@
Given a language $L\subseteq\Sigma^{*}$ its canonical right congruence $\sim_{L}$ is defined as follows: $x\sim_{L}y$ iff ${\forall z\in\Sigma^{*}}$ we have ${xz\in L}\iff{yz\in L}$. The Myhill-Nerode Theorem states that a language $L\subseteq\Sigma^{*}$ is regular iff $\sim_{L}$ is of finite index. Moreover, if $L$ is accepted by a complete DFA $\mathcal{A}$, then $\sim_{\mathcal{M}}$ refines $\sim_{L}$, where $\mathcal{M}$ is the automaton of $\mathcal{A}$. Finally, any complete DFA of minimum size that accepts $L$ has an automaton that is isomorphic to $\mathcal{M}_{\sim_{L}}$.
For an $\omega$-language $L\subseteq\Sigma^{\omega}$, its canonical right congruence $\sim_{L}$ is defined similarly, by quantifying over $\omega$-words. That is, $x\sim_{L}y$ iff ${\forall z\in\Sigma^{\omega}}$ we have ${xz\in L}\iff{yz\in L}$. If $L$ is a regular $\omega$-language then $\sim_{L}$ is of finite index, and for any complete DBA (resp., DCA, DPA, DRA, DSA, DMA) $\mathcal{A}$ that accepts $L$, $\sim_{\mathcal{M}}$ refines $\sim_{L}$, where $\mathcal{M}$ is the automaton of the acceptor.
@@
In light of the lack of a Myhill-Nerode result for regular $\omega$-languages, we define a restricted type of deterministic Büchi (resp., coBüchi, parity, Rabin, Streett, Muller) acceptor. Following the three-letter acronym notation introduced in the Acceptors subsection, for $X\in\{B,C,P,R,S,M\}$ we say that a DXA $\mathcal{A}$ recognizing language $L$ is fully informative if it is complete and its automaton is isomorphic to $\mathcal{M}_{\sim_{L}}$ (the automaton that corresponds to the canonical right congruence $\sim_{L}$ for the language, as defined above). A DXA is an IXA if it is fully informative. A DXA $\mathcal{B}$ is in $\mathbb{IXA}$ if there exists an IXA $\mathcal{A}$ such that ${\llbracket}\mathcal{B}{\rrbracket}={\llbracket}\mathcal{A}{\rrbracket}$. We note that every state of an IXA is reachable because every state of $\mathcal{M}_{\sim_{L}}$ is reachable. Since every state of a minimal automaton for a language $L$ in $\mathbb{IXA}$ for $\mathbb{X}\in\{\mathbb{B},\mathbb{C},\mathbb{P},\mathbb{R},\mathbb{S},\mathbb{M}\}$ correspond to an equivalence class of $\sim_{L}$ we refer to the $\mathbb{IXA}$ classes as the fully informative classes.
@@ 3.2. Computable Teachers and Learners
Under the assumptions (C1), (C2), and (C3), there is are a computable teacher $\mathbf{T}$ and a computable learner $\mathbf{L}$ such that $\mathbf{T}$ computes characteristic samples for $\mathbb{C}$ and $\mathbf{L}$.
@@
First we prove a lemma on using the majority vote algorithm with a finite set of concepts to construct a sample. If $\mathbb{D}$ is a finite nonempty set of concepts, we define the majority vote concept of $\mathbb{D}$, denoted $\mathcal{C}_{\textit{maj}}(\mathbb{D})$, to consist of all $x\in\mathcal{X}$ such that the cardinality of the set of concepts in $\mathbb{D}$ containing $x$ is at least as large as the cardinality of the set of concepts in $\mathbb{D}$ not containing $x$. The concept $\mathcal{C}_{\textit{maj}}(\mathbb{D})$ may or may not belong to $\mathbb{D}$ or $\mathbb{C}$. Given a finite nonempty set of concepts $\mathbb{D}$ and another concept $\mathcal{C}$, we define a sample, denoted $\textit{sample}(\mathbb{D},\mathcal{C})$ by executing the halving algorithm with $\mathbb{D}$ as the concept set and $\mathcal{C}$ as the target concept. In detail, initialize $\mathbb{D}_{0}=\mathbb{D}$, $i=0$, and $S_{0}=\emptyset$. While $\mathbb{D}_{i}$ is nonempty, compare $\mathcal{C}_{\textit{maj}}(\mathbb{D}_{i})$ and $\mathcal{C}$. If these are the same concept, then $S_{i}$ is output as the value of $\textit{sample}(\mathbb{D},\mathcal{C})$. Otherwise, let $x$ be the least element of $\mathcal{X}$ on which they differ, and add the pair $(x,b)$ to $S_{i}$ to get $S_{i+1}$, where $b=1$ if $x$ is in ${\llbracket}\mathcal{C}{\rrbracket}$ and $b=0$ otherwise. $\mathbb{D}_{i+1}$ is set to those concepts in $\mathbb{D}_{i}$ that are consistent with $(x,b)$, $i$ is set to $i+1$, and the while loop continues. If $\mathbb{D}_{i}$ is empty, then $S_{i}$ is the value of $\textit{sample}(\mathbb{D},\mathcal{C})$.
@@
The least $i$ such that $\mathbb{B}_{i}$ contains $\mathcal{C}_{k}$ is at most $\log_{2}\log_{2}(k)$ for $k>1$. So for $k>1$ by instantiating $i$ with $\log_{2}\log_{2}(k)$ in Eq. 3.1 we get that the cardinality of $\mathbf{T}(\mathcal{C})$ is at most $2\log_{2}(k)+\log\log(k)$. In the shortlex ordering of finite binary strings, the string $\mathcal{C}_{k}$ has length at least $\log_{2}k$, so the cardinality of $\mathbf{T}(\mathcal{C})$ is bounded by $O(|\mathcal{C}|)$.
@@
Let $\mathbb{C}$ be a concept class satisfying (C1), (C2), and (C3). There exist a computable teacher $\mathbf{T}$ and a computable learner $\mathbf{L}$ such that for every concept representation $\mathcal{C}$ in $\mathbb{C}$, $\mathbf{T}(\mathcal{C})$ is a characteristic sample for $\mathcal{C}$ and $\mathbf{L}$, and the cardinality of $\mathbf{T}(\mathcal{C})$ is bounded by $O(|\mathcal{C}|)$.

### v4

@@ 1. Introduction
One of the main obstacles in obtaining a polynomial learning algorithm for regular $\omega$-languages is that they do not in general have a Myhill-Nerode characterization; that is, there is no theorem correlating the states of a minimal automaton of some of the common automata types (Büchi, parity, Muller, etc.) to the equivalence classes of the right congruence of the language. The right congruence relation for an $\omega$-language $L$ relates two finite words $x$ and $y$ iff there is no infinite suffix $z$ differentiating them, that is $x\sim_{L}y$ (for $x,y\in\Sigma^{*}$) iff $\forall z\in\Sigma^{\omega}.\ xz\in L\iff yz\in L$. The quest for finding a polynomial query learning algorithm for a subclass of the regular $\omega$-languages, led to studying subclasses of languages for which such a relation holds. These languages are termed fully informative [AF18]. We use $\mathbb{IBA},\mathbb{ICA},\mathbb{IPA},\mathbb{IRA},\mathbb{ISA},\mathbb{IMA}$ to denote the classes of languages that are fully informative of type Büchi, coBüchi, parity, Rabin, Streett and Muller, respectively. A language $L$ is said to be fully informative of type $\mathbb{X}$ for $\mathbb{X}\in\{\mathbb{B},\mathbb{C},\mathbb{P},\mathbb{R},\mathbb{S},\mathbb{M}\}$ if there exists a deterministic automaton of type $\mathbb{X}$ that recognizes $L$ and is isomorphic to the automaton derived from $\sim_{L}$. While many properties of these classes are now known, in particular that they span the entire hierarchy of regular $\omega$-languages [Wag75], a polynomial learning algorithm for them is not known.
We show (in Sections 5-9) that the classes $\mathbb{IBA},\mathbb{ICA},\mathbb{IPA},\mathbb{IRA},\mathbb{ISA},\mathbb{IMA}$ can be identified in the limit using polynomial time and data. We further show (in Section 10) that there is a polynomial time algorithm to compute a characteristic sample given an acceptor $\mathcal{C}\in\mathbb{IXA}$. To show that these classes are also efficiently teachable we need polynomial time algorithms for inclusion and equivalence of automata of these types, that also return shortlex counterexamples in case of inequivalence. [^3] The formal definition of shortlex is deferred to Section 2. Such an algorithm is known to exist for the classes $\mathbb{NBA}$, $\mathbb{NCA}$, $\mathbb{NPA}$, since these classes have inclusion algorithms in NL [Sch10]. For the other classes a polynomial-time algorithm can be obtained following a reduction to model checking a certain fragment of CTL formulas [CDK93]. However this reduction does not yield shortlex counterexamples. We provide such algorithms in Sections 11-14.
@@ 2. Preliminaries
Let ${{\mathcal{M}}}_{1}$ and ${{\mathcal{M}}}_{2}$ be two deterministic complete automata with the same alphabet $\Sigma$, where for $i=1,2$, ${{\mathcal{M}}}_{i}=\langle\Sigma,Q_{i},(q_{\iota})_{i},\delta_{i}\rangle$. Their product automaton, denoted ${{\mathcal{M}}}_{1}\times{{\mathcal{M}}}_{2}$, is the deterministic complete automaton ${\mathcal{M}}=\langle\Sigma,Q,q_{\iota},\delta\rangle$ such that $Q=Q_{1}\times Q_{2}$ is the set of ordered pairs of states of ${{\mathcal{M}}}_{1}$ and ${{\mathcal{M}}}_{2}$; the initial state $q_{\iota}=((q_{\iota})_{1},(q_{\iota})_{2})$ is the pair of initial states of the two automata; and for all $(q_{1},q_{2})\in Q$ and $\sigma\in\Sigma$, $\delta((q_{1},q_{2}),\sigma)=(\delta_{1}(q_{1},\sigma),\delta_{2}(q_{2},\sigma))$. For $i=1,2$, let $\pi_{i}$ be projection onto the $i$-th coordinate, so that for a subset $S$ of $Q$, $\pi_{1}(S)=\{q_{1}\in Q_{1}\mid\exists q_{2}\in Q_{2}.\ (q_{1},q_{2})\in S\}$, and analogously for $\pi_{2}$.
@@
For finite words the acceptance condition is a set $F\subseteq Q$ and a run on a word $v$ is accepting if it ends in an accepting state, i.e., if $\delta(q_{\iota},v)$ contains an element of $F$. For infinite words, there are various acceptance conditions in the literature, and we consider six of them: Büchi, coBüchi, parity, Rabin, Streett and Muller, all based on the set of states visited infinitely often in a given run. For each model we define the related quantity of the size of the acceptor, taking into account the acceptance condition.
@@
We use three-letter acronyms for automata and classes concerning the common $\omega$-automata discussed above. The first letter is in $\{N,D,I\}$ and stands for non-deterministic, deterministic and isomorphic (or fully informative) which will be explained in the sequel. The second letter describes the acceptance condition, and the third letter $A$ stands for acceptor. That is, we use NBA, NCA, NPA, NRA, NSA, NMA (resp., DBA, DCA, DPA, DRA, DSA, DMA) for non-deterministic (resp., deterministic) Büchi, coBüchi, parity, Rabin, Street and Muller acceptors. We use blackboard font for the respective classes of representations. That is, we use $\mathbb{NBA}$, $\mathbb{NCA}$, $\mathbb{NPA}$, $\mathbb{NMA}$, $\mathbb{NRA}$ and $\mathbb{NSA}$ (resp., $\mathbb{DBA}$, $\mathbb{DCA}$, $\mathbb{DPA}$, $\mathbb{DRA}$, $\mathbb{DSA}$ and $\mathbb{DMA}$) for the corresponding class of representations. It is known that $\mathbb{NCA}$ and $\mathbb{DCA}$ recognize the same languages and that the classes $\mathbb{DCA}$ and $\mathbb{DBA}$ are distinct proper subclasses of the regular $\omega$-languages. The other classes are the full class of regular $\omega$-languages.
@@
    Let $\mathcal{B}=\langle\Sigma,Q,q_{\iota},\delta,F\rangle$, where $\mathcal{B}$ is an NBA. Define the NRA $\mathcal{R}=\langle\Sigma,Q,q_{\iota},\delta,\{(F,\emptyset)\}\rangle$. Then $\mathcal{B}$ and $\mathcal{R}$ are equivalent and have the same size. $\mathcal{R}$ is deterministic if $\mathcal{B}$ is.
@@
    Let $\mathcal{C}=\langle\Sigma,Q,q_{\iota},\delta,F\rangle$, where $\mathcal{C}$ is an NCA. Define the NSA $\mathcal{S}=\langle\Sigma,Q,q_{\iota},\delta,\{(F,\emptyset)\}\rangle$. Then $\mathcal{S}$ and $\mathcal{C}$ are equivalent and have the same size. $\mathcal{S}$ is deterministic if $\mathcal{C}$ is.
@@
Given a language $L\subseteq\Sigma^{*}$ its canonical right congruence $\sim_{L}$ is defined as follows: $x\sim_{L}y$ iff ${\forall z\in\Sigma^{*}}.\ {xz\in L}\iff{yz\in L}$. The Myhill-Nerode theorem states that a language $L\subseteq\Sigma^{*}$ is regular iff $\sim_{L}$ is of finite index. Moreover, if $L$ is accepted by a complete DFA $\mathcal{A}$, then $\sim_{\mathcal{M}}$ refines $\sim_{L}$, where $\mathcal{M}$ is the automaton of $\mathcal{A}$. Finally, any complete DFA of minimum size that accepts $L$ has an automaton that is isomorphic to $\mathcal{M}_{\sim_{L}}$.
For an $\omega$-language $L\subseteq\Sigma^{\omega}$, its canonical right congruence $\sim_{L}$ is defined similarly, by quantifying over $\omega$-words. That is, $x\sim_{L}y$ iff ${\forall z\in\Sigma^{\omega}}.\ {xz\in L}\iff{yz\in L}$. If $L$ is a regular $\omega$-language then $\sim_{L}$ is of finite index, and for any complete DBA (resp., DCA, DPA, DRA, DSA, DMA) $\mathcal{A}$ that accepts $L$, $\sim_{\mathcal{M}}$ refines $\sim_{L}$, where $\mathcal{M}$ is the automaton of the acceptor.
@@
In light of the lack of a Myhill-Nerode result for regular $\omega$-languages, we define a restricted type of deterministic Büchi (resp., coBüchi, parity, Rabin, Streett, Muller) acceptors. Following the three-letter acronym notation introduced in the Acceptors subsection, for $X\in\{B,C,P,R,S,M\}$ we say that a DXA $\mathcal{A}$ recognizing a language $L$ is fully informative if it is complete and its automaton is isomorphic to $\mathcal{M}_{\sim_{L}}$ (the automaton that corresponds to the canonical right congruence $\sim_{L}$ for the language, as defined above). A DXA is an IXA if it is fully informative. A DXA $\mathcal{B}$ is in $\mathbb{IXA}$ if there exists an IXA $\mathcal{A}$ such that ${\llbracket}\mathcal{B}{\rrbracket}={\llbracket}\mathcal{A}{\rrbracket}$. We note that every state of an IXA is reachable because every state of $\mathcal{M}_{\sim_{L}}$ is reachable. Since every state of a minimal automaton for a language $L$ in $\mathbb{IXA}$ for $\mathbb{X}\in\{\mathbb{B},\mathbb{C},\mathbb{P},\mathbb{R},\mathbb{S},\mathbb{M}\}$ corresponds to an equivalence class of $\sim_{L}$ we refer to the $\mathbb{IXA}$ classes as the fully informative classes.
@@ 3.2. Computable teachers and learners
Under the assumptions (C1), (C2) and (C3), there is are a computable teacher $\mathbf{T}$ and a computable learner $\mathbf{L}$ such that $\mathbf{T}$ computes characteristic samples for $\mathbb{C}$ and $\mathbf{L}$.
@@
First we prove a lemma on using the majority vote algorithm with a finite set of concepts to construct a sample. If $\mathbb{D}$ is a finite nonempty set of concepts, we define the majority vote concept of $\mathbb{D}$, denoted $\mathcal{C}_{\textit{maj}}(\mathbb{D})$, to consist of all $x\in\mathcal{X}$ such that the cardinality of the set of concepts in $\mathbb{D}$ containing $x$ is at least as large as the cardinality of the set of concepts in $\mathbb{D}$ not containing $x$. The concept $\mathcal{C}_{\textit{maj}}(\mathbb{D})$ may or may not belong to $\mathbb{D}$ or $\mathbb{C}$. Given a finite nonempty set of concepts $\mathbb{D}$ and another concept $\mathcal{C}$, we define a sample, denoted $\textit{sample}(\mathbb{D},\mathcal{C})$ by executing the halving algorithm with $\mathbb{D}$ as the concept set and $\mathcal{C}$ as the target concept. In detail, initialize $\mathbb{D}_{0}=\mathbb{D}$, $i=0$ and $S_{0}=\emptyset$. While $\mathbb{D}_{i}$ is nonempty, compare $\mathcal{C}_{\textit{maj}}(\mathbb{D}_{i})$ and $\mathcal{C}$. If these are the same concept, then $S_{i}$ is output as the value of $\textit{sample}(\mathbb{D},\mathcal{C})$. Otherwise, let $x$ be the least element of $\mathcal{X}$ on which they differ, and add the pair $(x,b)$ to $S_{i}$ to get $S_{i+1}$, where $b=1$ if $x$ is in ${\llbracket}\mathcal{C}{\rrbracket}$ and $b=0$ otherwise. $\mathbb{D}_{i+1}$ is set to those concepts in $\mathbb{D}_{i}$ that are consistent with $(x,b)$, $i$ is set to $i+1$, and the while loop continues. If $\mathbb{D}_{i}$ is empty, then $S_{i}$ is the value of $\textit{sample}(\mathbb{D},\mathcal{C})$.
@@
The least $i$ such that $\mathbb{B}_{i}$ contains $\mathcal{C}_{k}$ is at most $\log_{2}\log_{2}(k)$ for $k>1$. So for $k>1$ by instantiating $i$ with $\log_{2}\log_{2}(k)$ in Equation 3.1 we get that the cardinality of $\mathbf{T}(\mathcal{C})$ is at most $2\log_{2}(k)+\log\log(k)$. In the shortlex ordering of finite binary strings, the string $\mathcal{C}_{k}$ has length at least $\log_{2}k$, so the cardinality of $\mathbf{T}(\mathcal{C})$ is bounded by $O(|\mathcal{C}|)$.
@@
Let $\mathbb{C}$ be a concept class satisfying (C1), (C2) and (C3). There exist a computable teacher $\mathbf{T}$ and a computable learner $\mathbf{L}$ such that for every concept representation $\mathcal{C}$ in $\mathbb{C}$, $\mathbf{T}(\mathcal{C})$ is a characteristic sample for $\mathcal{C}$ and $\mathbf{L}$, and the cardinality of $\mathbf{T}(\mathcal{C})$ is bounded by $O(|\mathcal{C}|)$.